# SWE-Bench Pro: Can AI Agents Solve Long-Horizon Software Engineering Tasks?

### v1

@@ Abstract
We introduce SWE-Bench Pro, a substantially more challenging benchmark that builds upon the best practices of SWE-Bench [25], but is explicitly designed to capture realistic, complex, enterprise-level problems beyond the scope of SWE-Bench. SWE-Bench Pro contains 1,865 problems sourced from a diverse set of 41 actively maintained repositories spanning business applications, B2B services, and developer tools. The benchmark is partitioned into a public set with open access to problems sourced from 11 repositories, a held-out set of 12 repositories and a commercial set of 18 proprietary repositories where we have formal partnership agreements with early-stage startups. Problems in the held-out and the commercial set are not publicly accessible, but we release results on the commercial set. Our benchmark features long-horizon tasks that may require hours to days for a professional software engineer to complete, often involving patches across multiple files and substantial code modifications. All tasks are human-verified and augmented with sufficient context to ensure resolvability. In our evaluation of widely used coding models, under a unified scaffold, we observe that their performance on SWE-Bench Pro remains below 25% (Pass@1), with GPT-5 achieving the highest score to date at 23.3%. To better understand these limitations, we cluster the failure modes observed in the collected agent trajectories for a clearer characterization of the error patterns exhibited by current models. Overall, SWE-Bench Pro provides a contamination-resistant testbed that more faithfully captures the complexity and diversity of real-world software development, advancing the pursuit of truly autonomous software engineering agents at a professional level.
Figure 1: SWE-Bench Pro is a dataset with challenging, enterprise-level, long-horizon software engineering tasks. Frontier models, such as GPT-5 and Claude Opus 4.1, score less than 25% on SWE-Bench Pro with the SWE-Agent [22] scaffold. We design the dataset with contamination resistance, difficulty filtering, and human augmentation/verification.
@@ 1 Introduction
Large Language Model (LLM) agents have been widely adopted in modern software development workflows. SWE-bench [13] and related works [23, 24, 22, 25, 15] establish the task of issue resolution as a de-facto standard for assessing their capability and usefulness. In this setting, an agent is given an entire codebase, a task description (e.g., a bug report or feature request) in natural language and is instructed to produce a code patch that resolves the issue and passes the repository’s test suite. These benchmarks have been instrumental in demonstrating both the substantial potential and the persistent limitations of current models as SWE agents.
Notably, the state-of-the-art agents have reported over 70% pass rate on SWE-Bench-Verified [15], a subset of SWE-Bench that is verifiably solvable by human programmers. In the next 6 - 12 months, there will be diminishing feedback from SWE-Bench-Verified to improve coding agents. Towards this end, this paper is motivated to (1) mitigate existing issues in SWE-Bench and (2) generate high-quality coding problems for evaluating the progress of LLM agents after SWE-Bench is saturated. As a result, we introduce SWE-Bench Pro.
Current coding benchmarks face several limitations. First, many benchmarks are susceptible to contamination [21, 7, 26, 19], as exemplified by recent works [21, 7, 5] and social media posts [1, 25]. This risk arises because widely used open-source repositories—particularly those distributed under permissive licenses (e.g., MIT, Apache 2.0, BSD)—are prime candidates for inclusion in the large-scale web-crawled corpora used to pre-train LLMs [3]. As a result, constructing benchmarks from public GitHub repositories is inherently difficult, since many are already accessible as training data. Second, existing tasks may not adequately capture the complexity of real-world software engineering. For example, SWE-Bench Verified [13] includes a substantial proportion of relatively trivial problems (161 out of 500) that require only one- to two-line modifications. In contrast, industrial software engineering, particularly in enterprise settings, often demands multi-file modifications spanning hundreds of lines [9, 16]. This discrepancy raises concerns about whether current benchmarks truly reflect the challenges faced in practical development scenarios.
@@
Overall, LLM agents achieve only modest resolution rates on SWE-Bench Pro ($\leq$23.3% on the public set; $\leq$17.8% on the commercial set), substantially lower than the >>70% Pass@1 reported on SWE-Bench Verified [15]. We additionally observe a marked performance gap between the public and commercial sets, underscoring the greater complexity of enterprise codebases. Performance also varies systematically by programming language and repository: models generally perform better on Python and Go tasks, while several JavaScript/TypeScript repositories yield considerably lower results. To further characterize model behavior, we employ an LLM-as-a-judge analysis that surfaces distinct failure modes. Larger models (e.g., Opus 4.1) often fail on semantic or algorithmic correctness in large, multi-file edits, whereas smaller models (e.g., Qwen 3 32B) more frequently fail due to issues in syntax and formatting, tool use, or context management.
@@ 2 Related Work
The development of autonomous software engineering agents represents a convergence of advances in large language models, code generation benchmarks, and program synthesis techniques.
@@ 2.1 Code and Software Engineering Benchmarks
The evaluation of code generation capabilities has evolved from simple function-level tasks to complex repository-level challenges. Chen et al. [4] introduced HumanEval, a foundational benchmark of 164 handwritten programming problems that established the standard for measuring functional correctness in generated code. This was complemented by MBPP [2], which provided approximately 1,000 crowd-sourced Python problems designed for entry-level programmers. For more challenging algorithmic tasks, APPS [11] introduced 10,000 programming problems spanning from simple to complex algorithmic challenges.
The field has since recognized the limitations of function-level evaluation. Jimenez et al. [13] pioneered repository-level evaluation with SWE-bench, presenting 2,294 real GitHub issues from 12 Python repositories that require understanding entire codebases to resolve. This revealed a significant performance gap, with state-of-the-art models resolving only the simplest issues. Building on this foundation, Zan et al. [24] extended the approach to multiple programming languages with Multi-SWE-bench, covering Java, TypeScript, JavaScript, Go, Rust, C, and C++ with 1,632 expert-curated instances. Da et al. [6] shows that these instances can be used for RL training as well as evaluation.
Several benchmarks have focused on specific aspects of repository-level understanding. Ding et al. [8] introduced CrossCodeEval for cross-file code completion, requiring models to leverage context from multiple files within a repository. Liu et al. [14] developed RepoBench with three interconnected tasks specifically designed for evaluating repository-level auto-completion systems. More recently, Zhuo et al. [28] presented BigCodeBench, emphasizing code generation with diverse function calls and complex instructions. The emergence of multimodal challenges is exemplified by SWE-bench Multimodal [23], which extends evaluation to visual software domains. These benchmarks collectively demonstrate the increasing sophistication required for comprehensive evaluation of code generation systems. He et al. [10] explores the ability of languages models in optimizing code performance.
@@ 2.2 Software Engineering Agents
The development of autonomous agents capable of resolving real-world software engineering tasks has seen rapid progress. Yang et al. [22] introduced SWE-agent, emphasizing the critical importance of agent-computer interfaces (ACIs) in enabling effective code manipulation, achieving 12.5% resolution rate on SWE-bench. This work highlighted how interface design can be as important as model capabilities for agent performance. Zhang et al. [27] developed AutoCodeRover, which combines LLMs with sophisticated AST-based code search capabilities, achieving 19% on SWE-bench-lite while maintaining low operational costs.
The field has explored various architectural approaches to agent design. Wang et al. [18] presented OpenHands, an open platform supporting multiple agent types and coordination mechanisms, evaluated across 15 different benchmarks. Huang et al. [12] proposed AgentCoder, employing a multi-agent framework with specialized agents for programming, test design, and test execution, demonstrating the benefits of role specialization. Wang et al. [17] introduced CodeAct, which unified agent action spaces using executable Python code, showing performance improvements of up to 20% over JSON or text-based approaches. Interestingly, Xia et al. [20] challenged the complexity trend with Agentless, a simple localization-repair approach.
@@ 3 Dataset Overview
Figure 2: SWE-Bench Pro is designed to mimic real, challenging software engineering tasks – with larger changes, across multiple files, sourced from professional software engineering repositories. Frontier models, such as GPT-5 and Claude Opus 4.1, score >70% of SWE-Bench Verified but less than 25% on SWE-Bench Pro. Patches are generated with SWE-Agent [22] and evaluated on the public subset of SWE-Bench Pro.
@@ 3.1 Characteristics of SWE-Bench Pro
Industrially-Relevant, Diverse, and Challenging Tasks. First, all repositories selected in SWE-Bench Pro are actively maintained professional projects with substantial user bases, comprehensive documentation, and established development practices. In addition, we source commercial repositories. These repositories are private and sourced from startups, where we contacted the company and purchased their engineering repos. We sample repositories from a diverse range of topics, including consumer applications with complex UI logic, B2B platforms with intricate business rules, and developer tools with sophisticated APIs. Second, we limit each repository to contribute 50-100+ instances. This avoids the situation where models get an advantage by being especially good at a single repository, rewarding models that can truly generalize. Finally, we require edits to span multiple files and contain a substantial code change, similar to real software engineering tasks. Subsequently, SWE-Bench Pro problems are naturally challenging – the best model performance is around 25%.
@@ 3.2 Task Specification
Interface. A common false negative pattern in existing evaluation is that, while the interface is specified implicitly in the problem statement, models may misname classes or function names. Here, we explicitly define the class and function names expected by the tests to avoid the failure mode when relevant.
@@
Tests. Every task includes human-reviewed test suites with fail2pass tests that verify issue resolution and pass2pass tests that ensure existing functionalities remain intact. We first run the tests without the gold patch, then apply the gold patch to determine relevant test statuses. We notice that some tests can be dynamic or fail occasionally. To mitigate it, we run each set of tests 3 times and filter out any test that doesn’t pass consistently. Finally, we perform an additional round of verification on the fail2pass tests where we ask annotators to filter out tests which are too broad or not relevant to the task description.
@@ 4.1 Sourcing Problems
A valid problem requires a commit pair that satisfies two conditions. First, the instance commit must either fix a bug or introduce a feature. Second, the commit pair must include a test patch that verifies the correctness of the fix or feature through a fail2pass transition: applying the test patch to the base commit should cause test failures, while applying both the test patch and the gold patch should result in all tests passing.
Public repositories were selected to capture a representative spread of programming languages, project scales, and application domains. Repos are sourced based on several criteria, such as their similarity to professional programs, popularity, and their ability to extract end-to-end problems. Private repositories were sourced from Scale’s internal assets, including companies acquired through mergers and acquisitions, startups founded by Scale employees, and purchased codebases via external data partnerships. Unlike public repositories, these remain inaccessible to model developers, reducing the risk of data leakage and enforcing stricter generalization. They also mirror industrial-scale practices, with complex build systems, layered dependencies, and extensive testing frameworks, thereby presenting more demanding scenarios for SWE agents.
@@ 5 Results
OpenAI GPT-5 23.3
@@
Table 1: Model performance on the public set of SWE-Bench Pro (N=731). Models are evaluated using SWE-Agent [22], without any ambiguity (e.g. we provide the augmented problem statement, requirements, interface).
Model Resolve (%)
Claude Opus 4.1 17.8
OpenAI GPT-5 14.9
Gemini 2.5 Pro Preview 10.1
Claude Sonnet 4 9.1
OpenAI GPT-4o 3.6
Scaffold. We use the SWE-Agent [22] scaffold. We also explore another popular scaffold, Agentless [20]. However, we find that Agentless has difficulty in multi-file editing, thus, produces low evaluation scores. We focus on SWE-Agent for our results.
Evaluation settings. All models use the latest versions as of September 18th, 2025. For open-source LLMs, we use vllm to host each model. Models are hosted on a single node, with 8 H100 Nvidia GPUs. We enable tool-use when possible, for open-weight models, we use syntax parsing to enable tool-use. Models have a maximum of 200 turns. We use the same prompt for all models, which is a basic prompt outlining the task, format requirements for the agent and description for available tools.
Results. Table 2 shows the results of various models on SWE-Bench Pro. We report Pass@1 as the resolve rate. OpenAI GPT-5 and Claude Opus 4.1 achieve the highest resolve rates at 23.3% and 22.7% respectively, substantially outperforming smaller models. Claude 4 Sonnet also achieves a 16.3% resolve rate, while earlier generation models like DeepSeek Qwen-3 32B and OpenAI GPT-4o show considerably lower performance at 3.4% and 3.9% respectively. There is also a significant performance gap between the public and commercial set, where the best models score less than 20% in the commercial set, highlighting the difficulty of navigating enterprise codebases.
6 Analysis
In this section, we provide additional analysis for model performance on SWE-Bench Pro. We include analysis of performance on different types of issues, and failure modes of agent trajectories for different models.
6.1 Model Performance
6.2 Trajectory Failure Modes
We conduct an LLM-as-a-judge analysis for failure modes of different models, utilizing GPT-5 as the judge. Our work follows Yang et al. [22], who demonstrate 87% alignment of automated judgments with human categorization of failure modes.
Overall Submitted Not-Submitted Model Submitted Not-Submitted Wrong Syntax Incorrect Instruction Edge Other Tool-Use Long- Stuck Solution Error File Following Case Context in Loop Claude Opus 4.1 74.0% 26.0% 48.5% 32.7% 5.0% 2.6% 0.9% 10.3% 69.9% 26.8% 3.3% (681) (239) (330) (223) (34) (18) (6) (70) (167) (64) (8) GPT-5 36.9% 63.1% 51.7% 23.2% 5.6% 3.4% 0.4% 15.7% 97.6% 2.0% 0.4% (267) (457) (138) (62) (15) (9) (1) (42) (446) (9) (2) Claude Sonnet 4 42.2% 57.8% 23.7% 7.5% 3.4% 2.0% 0.0% 63.4% 8.9% 61.6% 29.5% (295) (404) (70) (22) (10) (6) (0) (187) (36) (249) (119) Gemini 2.5 53.7% 46.3% 33.6% 57.0% 4.7% 1.8% 0.0% 2.9% 84.0% 15.1% 0.9% Pro Preview (491) (424) (165) (280) (23) (9) (0) (14) (356) (64) (4) GPT-4o 72.1% 27.9% 45.2% 36.7% 11.2% 6.2% 0.0% 0.7% 100.0% 0.0% 0.0% (569) (220) (257) (209) (64) (35) (0) (4) (220) (0) (0) Qwen3 32B 48.7% 51.3% 24.4% 47.7% 21.2% 2.3% 0.0% 4.4% 86.0% 1.2% 12.8% (386) (406) (94) (184) (82) (9) (0) (17) (349) (5) (52)
7 Limitations and Future Work
In this section, we discuss limitations of our work and potential avenues for future work.
7.1 Limitations
Issue Scope. The current evaluation framework focuses primarily on issue resolution through code patches. Real-world software engineering encompasses broader activities such as system design, code review, documentation, and architectural decisions that are not captured in the current benchmark structure.
Reduction in Ambiguity. The human augmentation process, while improving problem clarity, may inadvertently make problems too prescriptive by providing excessive detail in requirements and interface specifications. In the real-world, problems are ambiguous, with potential follow-up or exploration needed to start the task.
7.2 Future Work
Expanded Language Coverage. Future iterations of SWE-Bench Pro should incorporate more diverse programming languages and frameworks to better represent the software development ecosystem. This includes languages like Java, C#, Rust, Kotlin, and emerging languages that may become prevalent in industry settings.
8 Conclusion
Acknowledgments
We would like to thank the contributors for their hard work on their dataset. Some of them are: Fernando Carabedo, Donnahue George Jr, Elías Muñiz. We are also deeply appreciative of the early-stage startups that partnered with us to provide proprietary commercial codebases, enabling a more realistic evaluation of AI agents in enterprise settings. Finally, we acknowledge the open-source communities behind the GPL-licensed repositories for their foundational work in software engineering, which inspired this benchmark. This research would not have been possible without these collective efforts.
Appendix
Appendix A Example Task Instance

### v2

@@ Abstract
We introduce SWE-Bench Pro, a substantially more challenging benchmark that builds upon the best practices of SWE-Bench [zhang2025swebench], but is explicitly designed to capture realistic, complex, enterprise-level problems beyond the scope of SWE-Bench. SWE-Bench Pro contains 1,865 problems sourced from a diverse set of 41 actively maintained repositories spanning business applications, B2B services, and developer tools. The benchmark is partitioned into a public set with open access to problems sourced from 11 repositories, a held-out set of 12 repositories and a commercial set of 18 proprietary repositories where we have formal partnership agreements with early-stage startups. Problems in the held-out and the commercial set are not publicly accessible, but we release results on the commercial set. Our benchmark features long-horizon tasks that may require hours to days for a professional software engineer to complete, often involving patches across multiple files and substantial code modifications. All tasks are human-verified and augmented with sufficient context to ensure resolvability. In our evaluation of widely used coding models, under a unified scaffold, we observe that their performance on SWE-Bench Pro remains below 45% (Pass@1). To better understand these limitations, we cluster the failure modes observed in the collected agent trajectories for a clearer characterization of the error patterns exhibited by current models. Overall, SWE-Bench Pro provides a contamination-resistant testbed that more faithfully captures the complexity and diversity of real-world software development, advancing the pursuit of truly autonomous software engineering agents at a professional level.
@@ 1 Introduction
Large Language Model (LLM) agents have been widely adopted in modern software development workflows. SWE-bench [jimenez2024swebench] and related works [yang2024swebenchmm, zan2024multiswe, yang2024sweagent, zhang2025swebench, swebenchverified] establish the task of issue resolution as a de-facto standard for assessing their capability and usefulness. In this setting, an agent is given an entire codebase, a task description (e.g., a bug report or feature request) in natural language and is instructed to produce a code patch that resolves the issue and passes the repository’s test suite. These benchmarks have been instrumental in demonstrating both the substantial potential and the persistent limitations of current models as SWE agents.
Current coding benchmarks face several limitations. First, many benchmarks are susceptible to contamination [xu2024benchmark, deng2024investigating, Zhang2024ACE, White2024LiveBenchAC], as exemplified by recent works [xu2024benchmark, deng2024investigating, cheng2025survey] and social media posts [aleithan2024swe, zhang2025swebench]. This risk arises because widely used open-source repositories—particularly those distributed under permissive licenses (e.g., MIT, Apache 2.0, BSD)—are prime candidates for inclusion in the large-scale web-crawled corpora used to pre-train LLMs [brown2020language]. As a result, constructing benchmarks from public GitHub repositories is inherently difficult, since many are already accessible as training data. Second, existing tasks may not adequately capture the complexity of real-world software engineering. For example, SWE-Bench Verified [jimenez2024swebench] includes a substantial proportion of relatively trivial problems (161 out of 500) that require only one- to two-line modifications. In contrast, industrial software engineering, particularly in enterprise settings, often demands multi-file modifications spanning hundreds of lines [hassan2009predicting, steidl2017evaluating]. This discrepancy raises concerns about whether current benchmarks truly reflect the challenges faced in practical development scenarios.
@@ 2.1 Code and Software Engineering Benchmarks
The evaluation of code generation capabilities has evolved from simple function-level tasks to complex repository-level challenges. chen2021evaluating introduced HumanEval, a foundational benchmark of 164 handwritten programming problems that established the standard for measuring functional correctness in generated code. This was complemented by MBPP [austin2021program], which provided approximately 1,000 crowd-sourced Python problems designed for entry-level programmers. For more challenging algorithmic tasks, APPS [hendrycks2021measuring] introduced 10,000 programming problems spanning from simple to complex algorithmic challenges.
The field has since recognized the limitations of function-level evaluation. jimenez2024swebench pioneered repository-level evaluation with SWE-bench, presenting 2,294 real GitHub issues from 12 Python repositories that require understanding entire codebases to resolve. This revealed a significant performance gap, with state-of-the-art models resolving only the simplest issues. Building on this foundation, zan2024multiswe extended the approach to multiple programming languages with Multi-SWE-bench, covering Java, TypeScript, JavaScript, Go, Rust, C, and C++ with 1,632 expert-curated instances. Da2025AgentRLVRTS shows that these instances can be used for RL training as well as evaluation.
@@ 2.2 Software Engineering Agents
The development of autonomous agents capable of resolving real-world software engineering tasks has seen rapid progress. yang2024sweagent introduced SWE-agent, emphasizing the critical importance of agent-computer interfaces (ACIs) in enabling effective code manipulation, achieving 12.5% resolution rate on SWE-bench. This work highlighted how interface design can be as important as model capabilities for agent performance. zhang2024autocoderover developed AutoCodeRover, which combines LLMs with sophisticated AST-based code search capabilities, achieving 19% on SWE-bench-lite while maintaining low operational costs.
@@ 3.1 Characteristics of SWE-Bench Pro
Industrially-Relevant, Diverse, and Challenging Tasks. First, all repositories selected in SWE-Bench Pro are actively maintained professional projects with substantial user bases, comprehensive documentation, and established development practices. In addition, we source commercial repositories. These repositories are private and sourced from startups, where we contacted the company and purchased their engineering repos. We sample repositories from a diverse range of topics, including consumer applications with complex UI logic, B2B platforms with intricate business rules, and developer tools with sophisticated APIs. Second, we limit each repository to contribute 50-100+ instances. This avoids the situation where models get an advantage by being especially good at a single repository, rewarding models that can truly generalize. Finally, we require edits to span multiple files and contain a substantial code change, similar to real software engineering tasks. Subsequently, SWE-Bench Pro problems are naturally challenging.
@@ 3.2 Task Specification
Interface. SWE-Bench Pro tasks include feature additions and code enhancements, such as refactoring, that involve creating or altering classes and methods. For these tasks, a common false negative in unit-test verifiers is when a model submits a valid solution with different interfaces than what the unit test is expecting. Here, we explicitly define the class and function names expected by the tests to avoid this failure mode when relevant.
@@
Tests. Every task includes human-reviewed test suites with fail2pass tests that verify issue resolution and pass2pass tests that ensure existing functionalities remain intact.
@@ 3.3 Public, Commercial, and Held-Out SWE-Bench Pro
Figure 1: Patches on SWE-Bench Pro are larger, more challenging, and require expert knowledge about a variety of topics. Patches are generated with SWE-Agent [yang2024sweagent] and evaluated on the public subset of SWE-Bench Pro.
@@ 5 Results
Claude Sonnet 4.5 43.6
Claude Sonnet 4 42.7
OpenAI GPT-5 (high) 41.8
Claude Haiku 4.5 39.5
Kimi K2 Instruct 27.7
OpenAI GPT-OSS 120B 16.2
Table 1: Model performance on the public set of SWE-Bench Pro (N=731). Models are evaluated using SWE-Agent [yang2024sweagent], without any ambiguity (e.g. we provide the augmented problem statement, requirements, interface).
Model Resolve (%)
Claude Opus 4.1 17.8
OpenAI GPT-5 (high) 15.7
OpenAI GPT-5 (medium) 14.9
Gemini 2.5 Pro Preview 10.1
Claude Sonnet 4 9.1
OpenAI GPT-4o 3.6
Scaffold. We use the SWE-Agent [yang2024sweagent] scaffold. We also explore another popular scaffold, Agentless [xia2024agentless]. However, we find that Agentless has difficulty in multi-file editing, thus, produces low evaluation scores. We focus on SWE-Agent for our results.
Evaluation settings. All models use the latest versions as of September 18th, 2025. For open-source LLMs, we use vllm to host each model. Models are hosted on a single node, with 8 H100 Nvidia GPUs. We enable tool-use when possible, for open-weight models, we use syntax parsing to enable tool-use. Models have a maximum of 50 turns. We use the same prompt for all models, which is the default prompt from [yang2024sweagent]. We use Opus 4.1 without extended thinking as reported by Anthropic in their own evaluation on SWE-Bench Verified.
Results. Table 2 shows the results of various models on SWE-Bench Pro. We report Pass@1 as the resolve rate. Claude Sonnet 4.5 and Claude Sonnet 4 achieve the highest resolve rates at 43.6% and 42.7% respectively, substantially outperforming smaller models. OpenAI GPT-5 (high) also achieves a 41.8% resolve rate, while Claude Haiku 4.5 demonstrates strong performance at 39.5%. Earlier generation models like Kimi K2 Instruct and OpenAI GPT-OSS 120B show considerably lower performance at 27.7% and 16.2% respectively. There is also a significant performance gap between the public and commercial set, where the best models score less than 20% in the commercial set, highlighting the difficulty of navigating enterprise codebases.
6 Analysis
In this section, we provide additional analysis for model performance on SWE-Bench Pro. We include analysis of performance on different types of issues, failure modes of agent trajectories for different models, and the efficacy of human augmentations for improving task resolvability. For compute constraints, analysis is done on trajectories with a max turn limit of 50 and max cost of \$2. Corresponding results are in Table 5.
Model
Problem Statement,
Requirements, Interface
Problem Statement
Only
OpenAI GPT-5 (high) 25.9% 8.40%
Claude Opus 4.1 22.7% 8.20%
Table 3: Comparison of model performance with and without human augmentations. The setting Problem Statement Only is without human augmentation. Without these augmentations, unit test verifiers are susceptible to false negatives.
6.1 Analysis on Model Performance on SWE-Bench Pro
Model behavior correlates with task complexity. The relationship between file count and resolve rate (Figure 3, top left) reveals distinct performance degradation patterns. Models maintain relatively stable performance for single-file problems but exhibit sharp declines as file count increases. Notably, the performance gap between frontier and smaller models widens dramatically beyond 3 files, with Claude Opus 4.1 and OpenAI GPT-5 maintaining above 10% resolve rates even for problems involving 10+ files, while open-source alternatives approach near-zero performance. This suggests that handling multi-file contexts requires capacity that current open-source models lack.
6.2 Ablation: Removing Human Augmentations
We perform an ablation in the importance of augmentations in SWE-Bench Pro, namely the requirements and interface. These augmentations provide the agent with enough information to mitigate false negatives from unit tests that expect specific APIs, signatures, or functional behavior. Table 3 shows results in two settings: our default setting (where we include the problem statement, requirements, and interface), and another setting where we include only the problem statement. Without the requirements and interface, both models tested (GPT-5 and Claude Opus 4.1) show significantly degraded performance. In this setting, agents are less constrained and can submit more diverse solutions (particularly for feature additions). Since unit tests expect a narrow set of solutions, verifiers are prone to false negatives, resulting in lower pass rates. Our human augmentations mitigate these false negatives by constraining the agent’s solution space such that verifiers are robust, providing appropriate context for measuring task resolution.
6.3 Trajectory Failure Modes
We conduct an LLM-as-a-judge analysis for failure modes of different models, utilizing GPT-5 as the judge. Our work follows yang2024sweagent, who demonstrate 87% alignment of automated judgments with human categorization of failure modes.
Categories. Here, we detail several of the common categories of failure modes. For the full category descriptions, view the Appendix. Wrong solution. The agent produces a syntactically valid patch that is functionally incorrect, incomplete, or fails to address the core problem. Tool-Use. Failure is attributed to the agent’s incorrect use of its available tools. This misuse prevents the agent from gathering necessary information or applying changes correctly. Syntax error. The agent successfully modifies the target files but introduces syntactic errors that render the codebase uncompilable or unrunnable. Incorrect file. This failure occurs when the agent correctly understands the high-level goal but fails to locate the correct source file or function for modification.
Overall Submitted Not-Submitted Model Submitted Not-Submitted Wrong Syntax Incorrect Instruction Edge Other Tool-Use Long- Stuck Solution Error File Following Case Context in Loop Claude Opus 4.1 74.2% 25.8% 50.3% 31.3% 4.9% 2.7% 0.8% 10.0% 68.0% 28.7% 3.4% (511) (178) (257) (160) (25) (14) (4) (51) (121) (51) (6) GPT-5 (high) 27.2% 72.8% 39.5% 29.3% 8.8% 4.8% 0.0% 17.7% 96.4% 2.5% 1.0% (147) (394) (58) (43) (13) (7) (0) (26) (380) (10) (4) Claude Sonnet 4 44.1% 55.9% 25.5% 7.7% 2.1% 2.1% 0.0% 62.6% 8.7% 57.4% 33.9% (235) (298) (60) (18) (5) (5) (0) (147) (26) (171) (101) Gemini 2.5 52.9% 47.1% 35.0% 56.5% 4.0% 1.9% 0.0% 2.7% 83.6% 14.3% 2.1% Pro Preview (377) (335) (132) (213) (15) (7) (0) (10) (280) (48) (7) GPT-4o 72.1% 27.9% 45.2% 36.7% 11.2% 6.2% 0.0% 0.7% 100.0% 0.0% 0.0% (569) (220) (257) (209) (64) (35) (0) (4) (220) (0) (0) Qwen3 32B 47.3% 52.7% 25.0% 48.7% 19.7% 1.6% 0.7% 4.3% 78.8% 1.5% 19.8% (304) (339) (76) (148) (60) (5) (2) (13) (267) (5) (67)
7 Limitations and Future Work
7.1 Limitations
7.2 Future Work
8 Conclusion
Appendix
Model Resolve (%)
OpenAI GPT-5 (high) 25.9
OpenAI GPT-5 (medium) 23.3
@@ 8 Conclusion
Table 5: Model performance on the public set of SWE-Bench Pro (N=731). Models are capped at 50 turns and a cost limit of \$2.
Appendix A Failure Mode Category Descriptions
  • Wrong solution. The agent produces a syntactically valid patch that is functionally incorrect, incomplete, or fails to address the core problem.
  • Tool-Use. Failure is attributed to the agent’s incorrect use of its available tools. This misuse prevents the agent from gathering necessary information or applying changes correctly.
  • Syntax error. The agent successfully modifies the target files but introduces syntactic errors that render the codebase uncompilable or unrunnable.
  • Incorrect file. This failure occurs when the agent correctly understands the high-level goal but fails to locate the correct source file or function for modification.
  • Endless File Reading:
    The agent enters a non-productive loop of exploratory actions, such as repeatedly reading the same files, searching for keywords, and viewing code snippets, without ever progressing to an implementation phase. It successfully gathers information but fails to synthesize it into a concrete code modification, eventually timing out or failing due to inaction.
  • Misunderstood Problem Statement:
    This category describes failures where the agent fundamentally misinterprets the task’s objective. Instead of implementing the required code changes, it pursues a tangential or incorrect goal, such as focusing on creating a complex runtime reproduction environment for a simple refactoring task. The agent’s actions are coherent with its flawed understanding but do not address the actual issue.
  • Other:
    A catch-all category for failures that do not fit into the more specific classifications above. This often includes trajectories where the agent exceeds computational or time limits (e.g., cost limits) due to an inefficient workflow, such as running an excessive number of verbose tests, rather than a single, clear technical mistake. It can also encompass a combination of minor, compounding issues.
Appendix B Example Task Instance